OpenStudy (anonymous):

[6.02] Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (btaylor):

solve for x: x=1+4y now put x into the second equation: 2(1+4y)-9y=3 2+8y-9y=3 -y = 1 y = -1 given your options, it must be D. But you would get x by plugging in -1 for y in the first equation: x-4(-1) = 1 x+4=1 x = -3